OpenStudy (anonymous):

the S.A of the top surface of the water in a circular swimming pool is about 206 ft estimate the radius of the pool to the nearest foot

OpenStudy (anonymous):

The surface area of the top of a cylindrical object is a circle. \[A = πr^{2}\]\[206 = πr^{2}\]JJust solve for r :)

hero (hero):

^What @Calcmathlete said

OpenStudy (anonymous):

what is "r"

hero (hero):

r = radius

OpenStudy (anonymous):

well yeah but is "r" 206?

hero (hero):

A = 206 as @Calcmathlete implied above

hero (hero):

We know that \(A = \pi r^2\) is the formula for Area of a circle. Replacing A with 206 in the formula gives \(206 = \pi r^2\) which implies that A = 206. The problem states: the surface area of the top surface of the water in a circular swimming pool is about 206 ft. In other words, \(A=206\).

hero (hero):

The only thing left to do is solve for r. You can use 3.14 for \(\pi\).

OpenStudy (anonymous):

thanks i passed my test @hero

hero (hero):

congrats